Please help me solve for x:
(9/x)+(9/x^2)+(9/x^3)+(109/x^4)=109.58

Answer by stanbon(75887)

solve for x:
(9/x)+(9/x^2)+(9/x^3)+(109/x^4)=109.58
Multiply thru by x^4 to get:
9x^3 + 9x^2 + 9x + 109 = 109.58x^4
---------------------
109.58x^4 - 9x^3-9x^2-9x-109 = 0
-----
Graph that to see solutions::
x = -0.978.. or x = 1.06.. etc
----------------
Cheers,
Stan H.